# CAPE: Connectivity-Aware Path Enforcement Loss for Curvilinear Structure Delineation

### v1

@@ Abstract
Promoting the connectivity of curvilinear structures, such as neuronal processes in biomedical scans and blood vessels in CT images, remains a key challenge in semantic segmentation. Traditional pixel-wise loss functions, including cross-entropy and Dice losses, often fail to capture high-level topological connectivity, resulting in topological mistakes in graphs obtained from prediction maps. In this paper, we propose CAPE (Connectivity-Aware Path Enforcement), a novel loss function designed to enforce connectivity in graphs obtained from segmentation maps by optimizing a graph connectivity metric. CAPE uses the graph representation of the ground truth to select node pairs and determine their corresponding paths within the predicted segmentation through a shortest-path algorithm. Using this, we penalize both disconnections and false positive connections, effectively promoting the model to preserve topological correctness. Experiments on 2D and 3D datasets, including neuron and blood vessel tracing demonstrate that CAPE significantly improves topology-aware metrics and outperforms state-of-the-art methods. The code will be made publicly available upon acceptance.
Keywords:
Segmentation Topology Optimization Biomedical Imaging Neuronal Process Tracing Blood Vessel Segmentation.
@@ 1 Introduction
In this paper, we propose CAPE (Connectivity-Aware Path Enforcement), a novel loss function that improves the connectivity of curvilinear structure reconstructions. Inspired by the Average Path Length Similarity (APLS) metric [5], CAPE evaluates connectivity by comparing entire paths instead of individual pixels. Specifically, it leverages Dijkstra’s algorithm to compute shortest paths in the pixel domain by mapping paths between nodes from the ground truth graph onto the predicted segmentation. This direct comparison enforces high-level topological correctness, leading to improved segmentation of curvilinear structures.
@@ 3 Methodology
Our approach is inspired by the Average Path Length Similarity (APLS) connectivity metric, which evaluates the connectivity of curvilinear structures but is inherently non-differentiable and unsuitable for direct integration into training segmentation networks. At a high level, our loss function measures the consistency of connectivity between the ground truth and the network’s prediction by comparing the costs of corresponding shortest paths. Specifically, we consider the ground truth graph $G=(V,E)$, where vertices $V$ represent key structural points and edges $E$ capture the true connections among foreground elements. Our network, $\mathbf{f}_{\theta}{}$, predicts a distance map $\hat{\mathbf{y}}$ in which each pixel indicates its distance from the nearest foreground structure (for brevity, we refer to image elements as pixels; in 3D data, these correspond to voxels, and the same methodology applies).
@@ 3.1 CAPE Loss
$\text{cost}(\text{path}_{\hat{\mathbf{y}}})=\sum_{n\in\text{path}_{\hat{\mathbf{y}}}}\hat{\mathbf{y}}(n)^{2},$
@@
$L_{\textsc{CAPE}}(\text{path}_{G},\text{path}_{\hat{\mathbf{y}}})=\left|\text{cost}(\text{path}_{G})-\text{cost}(\text{path}_{\hat{\mathbf{y}}})\right|=\text{cost}(\text{path}_{\hat{\mathbf{y}}}).$
@@
9:     $L_{\textsc{CAPE}}\leftarrow L_{\textsc{CAPE}}+cost(\text{path}_{\hat{\mathbf{y}}})$
@@
$\text{path}_{\hat{\mathbf{y}}}=\text{Dijkstra}(\hat{\mathbf{y}}\cdot M_{\text{dilated}},v^{\prime}_{1},v^{\prime}_{2}).$
@@ 3.2 Training Loss Integration
The total CAPE loss, $L_{\textsc{CAPE}}(G,\hat{\mathbf{y}})$, is defined as the sum of the losses computed for each ground truth path sampled from the graph $G$.
$L_{\textsc{CAPE}}(G,\hat{\mathbf{y}})=\sum_{\text{path}_{G}\in\mathcal{P}_{G}}L_{\textsc{CAPE}}(\text{path}_{G},\text{path}_{\hat{\mathbf{y}}})=\sum_{\text{path}_{G}\in\mathcal{P}_{G}}\text{cost}(\text{path}_{\hat{\mathbf{y}}}),$
@@
$L_{\textsc{total}}=L_{\text{MSE}}(\mathbf{y},\hat{\mathbf{y}})+\alpha L_{\text{CAPE}}(G,\hat{\mathbf{y}}),$
@@
Figure 3: Qualitative comparison of the test results in 2D datasets. The connectivity improves significantly when our approach is used.
@@ 4.3 Architectures and Baselines
For our 2D experiments, we use a 2D U-Net [22], featuring three down-samplings, two convolutional layers per level, max-pooling for the encoder, and bilinear up-sampling for the decoder. For 3D experiments, we employ a 3D U-Net with the same configuration. The 2D models are trained for 10k epochs and the 3D models are trained for 50k epochs, both with Adam optimizer [13] with a learning rate of $31e-$ 3 and weight decay of $31e-$ 3. We compare our results against MSE loss ($L_{MSE}$) between predictions and ground truths, a network trained with Perceptual loss using a pre-trained VGG19 [17, 24], clDice loss [23], and InvMALIS. For 2D datasets, we compare to inverse MALIS proposed in [19], and for 3D, we compare to the inverse MALIS computed on projections [20].

### v2

@@ Abstract
Promoting the connectivity of curvilinear structures, such as neuronal processes in biomedical scans and blood vessels in CT images, remains a key challenge in semantic segmentation. Traditional pixel-wise loss functions, including cross-entropy and Dice losses, often fail to capture high-level topological connectivity, resulting in topological mistakes in graphs obtained from prediction maps. In this paper, we propose CAPE (Connectivity-Aware Path Enforcement), a novel loss function designed to enforce connectivity in graphs obtained from segmentation maps by optimizing a graph connectivity metric. CAPE uses the graph representation of the ground truth to select node pairs and determine their corresponding paths within the predicted segmentation through a shortest-path algorithm. Using this, we penalize both disconnections and false positive connections, effectively promoting the model to preserve topological correctness. Experiments on 2D and 3D datasets, including neuron and blood vessel tracing demonstrate that CAPE significantly improves topology-aware metrics and outperforms state-of-the-art methods. Code is available at https://github.com/NeuraVisionLab/CAPE.
@@ 1 Introduction
In this paper, we propose CAPE (Connectivity-Aware Path Enforcement), a novel loss function that improves the connectivity of curvilinear structure reconstructions. Inspired by the Average Path Length Similarity (APLS) metric [5], CAPE evaluates connectivity by comparing entire paths instead of individual pixels. Specifically, it leverages Dijkstra’s algorithm to compute shortest paths in the pixel domain by mapping paths between nodes from the ground truth graph onto the predicted segmentation. This direct comparison enforces high-level topological correctness, leading to improved segmentation.
@@ 3 Methodology
Our approach is inspired by the APLS connectivity metric, which evaluates the connectivity of curvilinear structures but is inherently non-differentiable and unsuitable for direct integration into training segmentation networks. In APLS, edges are binary—either present or absent. In our differentiable adaptation, we replace these binary affinities with continuous ones (distance values predicted by our network). This allows path lengths to be computed as sums of network predictions, enabling optimization. At a high level, our loss function measures the consistency of connectivity between the ground truth and the network’s prediction by comparing the costs of corresponding shortest paths. Specifically, we consider the ground truth graph $G=(V,E)$, where vertices $V$ represent key structural points and edges $E$ capture the true connections among foreground elements. Our network, $\mathbf{f}_{\theta}{}$, predicts a distance map $\hat{\mathbf{y}}$ in which each pixel indicates its distance from the nearest foreground structure (for brevity, we refer to image elements as pixels; in 3D data, these correspond to voxels, and the same methodology applies). We predict distance maps instead of probability maps because distance map pixel values rise linearly as disconnections increase, more effectively penalizing longer or severe gaps. Probability maps, however, saturate near zero in disconnected regions, reducing loss sensitivity.
@@ 3.1 CAPE Loss
$\text{cost}(\text{path}_{\hat{\mathbf{y}}})=\sum_{n\in\text{path}_{\hat{\mathbf{y}}}}\hat{\mathbf{y}}(n)^{2},$ (1)
@@
[…]
@@
[…]
@@
$\text{path}_{\hat{\mathbf{y}}}=\text{Dijkstra}(\hat{\mathbf{y}}\cdot M_{\text{dilated}},v^{\prime}_{1},v^{\prime}_{2}).$ (3)
@@ 3.2 Training Loss Integration
[…]
@@
[…]
@@
Brain Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Qualitative comparison of the test results in 2D and 3D datasets. The arrows highlight regions that remain disconnected in alternative methods but are successfully connected using CAPE. The connectivity improves significantly when our approach is used.
@@ 4.3 Architectures and Baselines
For our 2D experiments, we use a 2D U-Net [21], featuring three down-samplings, two convolutional layers per level, max-pooling for the encoder, and bilinear up-sampling for the decoder. For 3D experiments, we employ a 3D U-Net with the same configuration. The 2D models are trained for 10k epochs and the 3D models are trained for 50k epochs, both with Adam optimizer [12] with a learning rate of 1e31e-3 and weight decay of 1e31e-3. We compare our results against MSE loss ($L_{MSE}$) between predictions and ground truths, a network trained with Perceptual loss using a pre-trained VGG19 [16, 23], clDice loss [22], and InvMALIS. For 2D datasets, we compare to inverse MALIS proposed in [18], and for 3D, we compare to the inverse MALIS computed on projections [19].
@@ 5 Conclusion and Future Work
[…]
The authors have no competing interests to declare that are relevant to the content of this article.